OpenStudy (anonymous):

A small publishing company is planning to publish a new book. The production costs will include one-time fixed costs (such as editing) and variable costs (such as printing). There are two production methods it could use. With one method, the one-time fixed costs will total $49,094 , and the variable costs will be $12.25per book. With the other method, the one-time fixed costs will total $16,334 and the variable costs will be $21 per book. For how many books produced will the costs from the two methods be the same?

OpenStudy (anonymous):

@texaschic101 do you know the answer?

OpenStudy (texaschic101):

49,094 + 12.25b = 16,334 + 21b 49,094 - 16,334 = 21b - 12.25b 32760 = 8.5b 32,760/8.5 = b 3854 = b check... 49094 + 12.25(3854) = 16,334 + 21(3854) 49094 + 47211.50 = 16,334 + 80,934 96305.50 = 97268 ? I can't seem to get it to equal.....???

OpenStudy (anonymous):

maybe the answer is just 0?

OpenStudy (texaschic101):

@genius12 ....can you help